OpenStudy (anonymous):

Another question I need help with.... What is the distance between P(–4, 3) and Q(6, 1)? Round to the nearest tenth.

OpenStudy (anonymous):

sqrt((6-(-4))^2+(3-1)^2))=sqrt(100+4)=sqrt(104)=10.1=10

OpenStudy (lalaly):

\[\huge{d=\sqrt{(x_2-x_1)^2+(y_2-y_1)^2}}\] \[\huge{d=\sqrt{(10)^2+(-2)^2}} = \sqrt{104}\]

OpenStudy (anonymous):

Thank you both!